# The Transition from Giant Planets to Brown Dwarfs beyond 1 au from the Stellar Metallicity Distribution

### v1

@@ 4.4 Implications for Planet Formation Theory
If we assume low values of $\alpha$ and typical disk depletion rates, high initial disk masses are required to form the most massive planets. Tanaka et al. (2020) predicted that planets with masses up to $\sim 20\,M_{\rm Jup}$ can form in disks with $\alpha=10^{-3}$ and initial masses of $\sim 0.1\,M_{\star}$ (roughly $10\times$ higher than the minimum mass solar nebula, assuming a Sun-like host star). Whether disks with these masses actually exist is still unknown, but the fact that companions with masses of $10-20\,M_{\rm Jup}$ are so rare means that high initial masses are only required in a small fraction of disks to reproduce the observed planetary mass function (see J. Van Zandt et al. in review for more info). In other words, a maximum planet mass of $\sim 27\,M_{\rm Jup}$ is consistent with the notion that $10-20\,M_{\rm Jup}$ planets represent the tail end of planet formation, where formation efficiency is low due to the small fraction of disks with the appropriate initial conditions.
@@ 5 Conclusions
Using a hierarchical Bayesian model, we inspected the distribution of giant planets, brown dwarfs, and low-mass stellar companions with orbital separations beyond 1 au detected by the California Legacy Survey (Rosenthal et al., 2021) in the companion mass – stellar metallicity plane. We found that the sample is well-described by a change-point model, in which two distinct stellar metallicity distributions are separated by a transition mass, $\gamma$. We calculated the transition mass to be $\gamma=27_{-8}^{+12}\,M_{\rm Jup}$, with relatively low-mass companions preferring to orbit relatively metal-rich stars ($\rm{[Fe/H]}=0.17\pm 0.12$ dex) and relatively high-mass companions orbiting relatively metal-poor stars ($\rm{[Fe/H]}=-0.03\pm 0.10$ dex). In addition, we found that the low-mass population has a stronger preference for low-eccentricity orbits (see G. Gilbert et al. 2025 for a more careful analysis of eccentricities). Overall, these findings are consistent with predictions of substellar companion formation: those with relatively low masses form bottom up, whereas those with relatively high-masses for top down. Our estimate of $\gamma$ broadly agrees with previous estimates based on other metrics.
@@
The authors thank Heather Knutson, Howard Isaacson, and Benjamin J. Fulton for their constructive conversations and feedback. SG is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303922.

### v2

@@ 4.4 Implications for Planet Formation Theory
If we assume low values of $\alpha$ and typical disk depletion rates, high initial disk masses are required to form the most massive planets. Tanaka et al. (2020) predicted that planets with masses up to $\sim 20\,M_{\rm Jup}$ can form in disks with $\alpha=10^{-3}$ and initial masses of $\sim 0.1\,M_{\star}$ (roughly $10\times$ higher than the minimum mass solar nebula, assuming a Sun-like host star). There is tentative evidence that some disks can reach these initial masses (Speedie et al., 2024, 2025), and the fact that companions with masses of $10-20\,M_{\rm Jup}$ are so rare means that high initial masses are only required in a small fraction of disks to reproduce the observed planetary mass function (see Van Zandt et al. 2025 for more info). In other words, a maximum planet mass of $\sim 27\,M_{\rm Jup}$ is consistent with the notion that $10-20\,M_{\rm Jup}$ planets represent the tail end of planet formation, where formation efficiency is low due to the small fraction of disks with the appropriate initial conditions.
@@ 5 Conclusions
Using a hierarchical Bayesian model, we inspected the distribution of giant planets, brown dwarfs, and low-mass stellar companions with orbital separations beyond 1 au detected by the California Legacy Survey (Rosenthal et al., 2021) in the companion mass – stellar metallicity plane. We found that the sample is well-described by a change-point model, in which two distinct stellar metallicity distributions are separated by a transition mass, $\gamma$. We calculated the transition mass to be $\gamma=27_{-8}^{+12}\,M_{\rm Jup}$, with relatively low-mass companions preferring to orbit relatively metal-rich stars ($\rm{[Fe/H]}=0.17\pm 0.12$ dex) and relatively high-mass companions orbiting relatively metal-poor stars ($\rm{[Fe/H]}=-0.03\pm 0.10$ dex). In addition, we found that the low-mass population has a stronger preference for low-eccentricity orbits (see Gilbert et al. 2025 for a more careful analysis of eccentricities). Overall, these findings are consistent with predictions of substellar companion formation: those with relatively low masses form bottom up, whereas those with relatively high-masses for top down. Our estimate of $\gamma$ broadly agrees with previous estimates based on other metrics.
@@
We thank the anonymous referee for their comments that significantly strengthened this paper. We also thank Heather Knutson, Howard Isaacson, and Benjamin J. Fulton for their constructive conversations and feedback. SG is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303922.